OpenStudy (anonymous):

\sqrt{2x+3}=2+\sqrt{x-2}

OpenStudy (hba):

\[2+\sqrt{x-2}\] Is it equal to 0 ?

OpenStudy (anonymous):

Yeah

OpenStudy (hba):

\[2+\sqrt{x-2}=0\] 1) Subtract 2 on both sides 2) Squaring both sides

OpenStudy (hba):

Start working @getexcool

OpenStudy (anonymous):

i got 6??

OpenStudy (hba):

@getexcool Respond....

OpenStudy (anonymous):

Well the question was suppose to be Sqrt:2x+3=2+sqrt:x-2

OpenStudy (anonymous):

For the first part I got 2x+3 cuz I am suppose to square both sides with 2

OpenStudy (hba):

The question is \[\sqrt{2x+3}=2+\sqrt{x-2}\] ?

OpenStudy (anonymous):

That's correct

OpenStudy (hba):

Apply squaring both sides on this \[( \sqrt{2x+3})^2=(2+\ \sqrt{x-2})^2\]

OpenStudy (hba):

Use (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

I got x=3 is that the right answer

OpenStudy (anonymous):

@hba

OpenStudy (hba):

Put the value in your question and check if LHS=RHS if yes then you are correct :)